OpenStudy (daniellelovee):

Use the correlation coefficient to find the coefficient of determination. What percent of variation in the response variable can be explained by changes in the explanatory variable? r = -0.5

OpenStudy (daniellelovee):

coefficient of determination: 0.1; 10% of the variation in the response variable can be explained by changes in the explanatory variable. coefficient of determination: 0.25; 25% of the variation in the response variable can be explained by changes in the explanatory variable. coefficient of determination: 0.5; 50% of the variation in the response variable can be explained by changes in the explanatory variable. coefficient of determination: 0.71; 71% of the variation in the response variable can be explained by changes in the explanatory variable.

OpenStudy (daniellelovee):

@AlexandervonHumboldt2

OpenStudy (daniellelovee):

@tkhunny

OpenStudy (tkhunny):

What do your course materials say?

OpenStudy (tkhunny):

Don't make me read it. I already know what it says. What are the salient points?

OpenStudy (tkhunny):

Which is the Coeffi8cient of Variation? r, r^2, or something else?

OpenStudy (daniellelovee):

idk im sorry I have no idea whatsoever how to solve this thats why I would like your help

OpenStudy (tkhunny):

The REAL help is you learning how to learn. Sometimes it takes searching. Sometimes it takes explanations. Sometimes it takes examples. I like to vary it.

OpenStudy (daniellelovee):

can you explain? or do what you think best please I really want and need to learn it

OpenStudy (tkhunny):

If the Coefficient of Determination explains a percentage of the variation, it can't very well be negative. Thus, it must be r^2, not r.

OpenStudy (daniellelovee):

so -0.5^2?

OpenStudy (daniellelovee):

1/4

OpenStudy (daniellelovee):

?

OpenStudy (tkhunny):

Careful with notation. -0.5^2 might mean -(0.5^2) = -0.25 (-0.5)^2 = 0.25 and cannot be mistaken for the negative value.

OpenStudy (daniellelovee):

yeah I was wondering that but I thought it was the first one

OpenStudy (tkhunny):

In some computer languages, -0.5^2 = (-0.5)^2 = 0.25 In others, it is as shown above. You have to know or you will not get your mars lander in the right orbit. :-)

OpenStudy (tkhunny):

Adding parentheses for clarification is almost never a mistake.

OpenStudy (daniellelovee):

yeah I learned that some time ago lol but was the positive answer the correct one or not?

OpenStudy (daniellelovee):

@satellite73

OpenStudy (tkhunny):

Review: If the Coefficient of Determination explains a percentage of the variation, it can't very well be negative. Thus, it must be r^2, not r. r^2 cannot be negative.

OpenStudy (daniellelovee):

ohh ok got it therefore my first answer was correct, im ready for the next step :)

OpenStudy (tkhunny):

In 5-8, you are given r. -1 <= r <= +1 and 0 <= r^2 <= 1

OpenStudy (tkhunny):

Yea!!

OpenStudy (daniellelovee):

alright does that mean tha B is the correct choice?

OpenStudy (tkhunny):

I don't see a "B". I see only numbered problems.

OpenStudy (daniellelovee):

coefficient of determination: 0.25; 25% of the variation in the response variable can be explained by changes in the explanatory variable.

OpenStudy (daniellelovee):

is that it?

OpenStudy (tkhunny):

Wow! I forgot the problem started all the way up there. Yes!!

OpenStudy (daniellelovee):

alright thank you for you help :) and for teaching me